Problem 2 EM Industries is considering a new project involving the acquisition of a new machine that would replace an older machine currently in use. The new machine costs $750,000 (at t-0) and can be sold at the end of its expected 4-year operating life for $100,000 (at t-4). The new machine takes up more space and EM will need to move maintenance and cleaning supplies that used to be stored next to the machine to a small storage room that could otherwise be sublet for $25,000 a year (at tl to t-4). The old machine was bought 3 years ago for $700,000 and can be sold for S150,000 today or for $50,000 in 4 years. Both, the old and the new machine belong to asset class 43 with a CCA rate of 30%. Management believes that the company will have other class 43 assets in four years when the equipment would be sold. EM Industries has just paid $55,000 for a study which indicates that the new machine will reduce annual manufacturing expenses by $200,000 per year (at t- to t4. Since the new machine is more reliable, the plant will need to keep fewer spare parts in stock. Management expects that inventory levels can be reduced by S35,000 (at t-0) when the new machine is installed (note, at the end of the project, this change in net working capital will be reversed, i.e., inventory levels will increase again by S35,000 at the end of year 4). EMs marginal tax rate is 35%, and its required rate of return (RRR) is 13% a) What is the initial cash outlay (the total cash flow at t-0)? b) What is the second years cash flow (excluding the CCA Tax Shield)? c) What is the last years cash flow (excluding the CCA Tax Shield)? d) What is the year 3 CCA? e) What is the PV CCA Tax Shield? f) What is the NPV of the replacement project?
Answer #1

STATEMENT OF CASHFLOWS (EXCLUDING CCA TAX SHIELD)

PARTICULARS                  YEAR 0            YEAR1        YEAR2            YEAR3           YEAR4

1. cost of new asset           (750000)

2. salvage value of old asset 150000

3. decrease in working capital   35000

4. saving in costs                                     200000        200000       200000     200000

5.less: opportunity costs                                      25000        25000            25000        25000

6. net savings in costs                                         175000         175000       175000      175000

7. after tax savings in costs(col 6 x 65%)            113750          113750       113750      113750

8.net salvage value of new machine                                                                              100000

9. increase in working capital                                                                                         (35000)

10.initial cash outlay                         (565000)

(col 1 + 2 + 3)

11. operating cashflow                                       113750        113750    113750          113750

(excluding CCa tax shield)

12. cashflow year4       

( col 8 + 9 +11)                                   178750                                                       

WORKING NOTES TO A\BOVE TABLE

1.loss of sublet charges of 25000 is treated as opportunity cost

2.cost of study of 55000 is SUNK cost . It will remain the same whether the project is accepted or not. So not considered in cashflow statement

3. tax rate is 35%. So after tax cost is 65%

ANSWER A

initial cash outlay = -565000 as per table above

ANSWER B

2nd year cashflow = 113750 as per table

ANSWER C

178750 as per table

ANSWER D

CALCULATION OF CCA

YEAR     BEGINING UCC             CCA                    END OF YEAR UCC

1             750000                      112500                      637500

2             637500                     191250                     446250

3             446250                      133875                     312375

4             312375                         93712.5               218662.5

In the above table , the opening balance of equipment is 750000. In the first year the half year rule applies.So we calculate CCA on half of 750000 that is 750000 x 1/2 x 30%=   112500. This leaves a balance of 750000-112500= 637500. So in tth next year CCA is 30 % of 637500 and so on.

as per the above table year 3 CCA is 133875

ANSWER E

PV CCA TAX SHIELD

PV of CCA tax shield is given by the following formula

\frac{C d T}{r + d}    X \frac{(1+0.5r)}{(1+r)}       -    \frac{(Sd T)}{(r + d)} X \frac{1}{(1+r)^{t}}

where C is initial cost of asset = 750000

d is CCA rate = 0.30

T is firm's tax rate = 0.35

r is discount rate = 0.13

S is salvage value of asset = 100000

Using this formula , PV=

\frac{750000*.3 *.35}{(0.13 +0.30)}   X \frac{1 + 0.5 * 0.13}{(1 +0.13 )}     - \frac{100000 *0.3 * 0.35}{(0.13 +0.30 )}    X \frac{1}{(1 + 0.13)^{4}}

= 78750/0.43 X 1.065/1.13   - (10500/0.43) (1/16.31)

183139.53 X 0.94 - (24418.60 X 0.06)

172151.16 - 1465.12=170686.04

answerf

NPV of the replacement project

NPV of total cashflows excluding CCA tax shields( as given in first ta\ble) + present value of CCA tax shields as calculated above